# Connection between gravity and quantum mechanics: an algebraic approach to cosmological issues

### v1

@@ XII Discussion and Conclusion
Our article examines PBHs as crucial factors in the universe’s evolution, specifically the early universe–relating to the Planck scale physics, particularly in relation to quantized gravity and cosmic events. The suggested model links the quantization of black hole characteristics with fundamental physics and cosmology, providing insights into high-energy astrophysics, cosmic structure development. This concept achieves the unification of quantum physics and gravity by the groundbreaking discovery of condensed light, namely, the self-gravitating photon Bose-Einstein condensate. This technique is fundamentally based on a geometric understanding that connects the Compton wavelength of the quantum photon condensate to the gravitational radius of PBHs. This connection elucidates a range of quantized attributes—encompassing geometrical, gravitational, thermodynamic, informational, and radiative properties—illuminating the fundamental principles that dictate the formation, development, and ultimate extinction of successive generations of PBHs.
@@
This advancement unveiled new aspects of Planck-scale and primordial black hole physics, facilitating the creation of efficient and robust mathematical instruments. These tools have facilitated accurate computations and resolutions to various fundamental issues in physics and cosmology, encompassing the beginning of time, Big Bang cosmology, the genesis and development of dark matter and dark energy, as well as the mystery of baryogenesis and baryonic asymmetry. The model also examines the hyper-exponential growth of SMBHs, the birth of primordial galaxies, the evolution of large-scale structures, and the Universe’s two-phase accelerated expansion. Furthermore, it offers very plausible solutions for persistent problems such as the horizon problem, the cosmological constant problem, the information loss conundrum, and the core-cusp dilemma.
The most striking aspect of our findings is that they almost align remarkably well with the predictions and observations of $\Lambda$CDM cosmology. This agreement is particularly notable because our results were derived independently, without relying on the standard equations and assumptions typically associated with $\Lambda$CDM.
It should be noted that, in our article, a method based on the membrane formalism of Price and Thorne [39], through an algebraic technique, is used to show signs of inhomogeneity caused by quantum fluctuations at the Planck scale. These results may fit with recent progress in inhomogeneous cosmology, as pointed out by Sarkar [111], who looked at models that are different from the homogeneous FLRW framework in great detail. Specifically, Sarkar’s work underscores the need for alternative mechanisms of structure formation, particularly in inhomogeneous and anisotropic universes, as discussed by Krasinski [112]. While inhomogeneous models have not yet been as thoroughly explored as FLRW models in terms of structure formation, empirical data increasingly supports a non-uniform cosmic landscape [113, 114, 115, 116]. These observations collectively indicate that large-scale inhomogeneities may play a significant role in the evolution of cosmic structures, suggesting a potential re-evaluation of our standard cosmological models. This article does not address inhomogeneity in detail, as it falls outside the scope of this study.
This research explores the relationship between quantum mechanics and cosmology by analyzing PBHs in terms of their quantum properties and making comparisons with BECs. The founding relationship can be experimentally investigated within the framework of analog gravity, building on the foundational work of Belgiorno et al. [33] and Agullo et al. [34]. The ultimate goal is to develop a unified theory integrating quantum mechanics and general relativity, paving the way for future theoretical advancements and space exploration.
Acknowledgement: The authors extend their gratitude to Prof. R. Ruffini, Prof. G. Vereshchagin, Prof. V. Mukhanov, Prof. S. Ansoldi, Prof. A. Pitteelli, Prof. A. Gaina, Prof. V. Galamaga, Prof. P. Panchadhyayee, I. Borşevici, and S. Borşevici for their valuable and insightful discussions, which greatly contributed to the development of this work.

### v2

@@ XII Conclusion
Our article examines PBHs as crucial factors in the universe’s evolution, specifically the early universe–relating to the Planck scale physics, particularly in relation to quantized gravity and cosmic events. The suggested model links the quantization of black hole characteristics with fundamental physics and cosmology, providing insights into high-energy astrophysics, and cosmic structure formation. This concept achieves the unification of quantum physics and gravity by the groundbreaking discovery of condensed light, namely, the self-gravitating photon Bose-Einstein condensate. This technique is fundamentally based on a geometric understanding that connects the Compton wavelength of the photon condensate to the gravitational radius of PBHs. This connection elucidates a range of quantized attributes—encompassing geometrical, gravitational, thermodynamic, informational, and radiative properties—illuminating the fundamental principles that dictate the formation, development, and ultimate extinction of successive generations of PBHs.
@@
This advancement unveiled new aspects of Planck-scale and PBH physics, facilitating the creation of efficient and robust mathematical instruments. These tools have facilitated accurate computations and resolutions to various fundamental issues in physics and cosmology, encompassing the beginning of time, Big Bang cosmology, the genesis and development of dark matter and dark energy, as well as the mystery of baryogenesis and baryonic asymmetry. Our model also examines the hyper-exponential growth of SMBHs, the birth of primordial galaxies, the evolution of large-scale structures, and the Universe’s two-phase accelerated expansion. Furthermore, it offers very plausible solutions for persistent problems such as the horizon problem, the cosmological constant problem, the information loss conundrum, and the core-cusp dilemma.
The most notable aspect of our results is their remarkable alignment with the predictions and observations of $\Lambda$CDM cosmology, which has been supported by many observations, including data from the JWST. This agreement is especially significant since our findings were obtained independently, without dependence on the conventional equations and assumptions often linked to $\Lambda$CDM.
The proposed framework provides a systematic method for analyzing quantum-gravitational processes at the Planck scale and addresses specific cosmological phenomena. However, its assumptions and predictions require further experimental and observational confirmation and comparison with other models. The main objective of this work is to provide the principal direction for further theoretical investigations and space exploration.
Acknowledgement: The authors extend their gratitude to Prof. R. Ruffini, Prof. G. Vereshchagin, Prof. V. Mukhanov, Prof. E. Guendelman, Prof. S. Ansoldi, Prof. A. Pitteelli, Prof. A. Gaina, Prof. V. Galamaga, Prof. P. Panchadhyayee, I. Borşevici, and S. Borşevici for their valuable and insightful discussions, which greatly contributed to the development of this work. V.B. extends his gratitude to the organizing committee and session chairpersons of the parallel session titled ”Inflation: Perturbations, Initial Singularities, and Emergent Universes” at the 17th Marcel Grossmann Meeting in Pescara, Italy, for the opportunity to present his talk at the event. This paper is the comprehensive and thoroughly revised version of the MG17 presentation. G.M. would like to extend thanks to all the undergraduate, postgraduate, and doctoral students who significantly enriched him.
@@
Declaration of competing interest: The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of generative AI in scientific writing: The authors state that they do not support the use of AI tools to analyze and extract insights from data as part of the study process.